f(x)=5x2+4x
(a)f(3)
(b)f(2a)

Answer by checkley75(3666)

F(X)=5X^2+4X
F(3)=5*3^2+4*3
F(3)=5*9+12
F(3)=45+12
F(3)=57 ANSWER.
--------------------------------------------------
F(2A)=5(2A)^2+4*2A
F(2A=5*4A^2+8A
F(2A)=20A^2+8A
F(2A)=4A(5A+2)